### v1

# Estimating the D-line Energy Splitting of Alkali Metals Using a Modified Perturbation Method

Abstract
[…]
Keywords: spin-orbit effect; sodium doublet; rubidium doublet; perturbation theory; screening constant
1 Introduction
The principle line spectra of an excited hydrogen gas sample was described empirically by J. J. Balmer in 1885. However, a splitting in some of these lines, or so called fine structure, was soon found to exist. About a decade later, Michelson and Morley discovered fine structure splitting in the spectrum of sodium [1]. By ${}^{\text{th}}$ century, the Balmer line spectra of hydrogen had been described theoretically by the Bohr formula where lines were taken to be the result of energy transitions between quantum states with principle quantum number $n=2$ and $n=3,4,5,\cdots$.
After the appearance of Bohr’s formula, a quantitative explanation for fine structure in hydrogen spectra would soon be developed. Since that time a variety of complex splittings have been observed in excited gas spectra but we will restrict our discussion here to a particular class of fine structure. This type of splitting occurs when one principle line is separated into two closely spaced bands, so called doublets, or D lines.
After the discovery of electron spin, a model for coupling between spin and orbital angular momentum was developed to explain the origin of the doublet splitting in hydrogen. This theory was further refined in 1927 by L. H. Thomas who included relativistic effects thus leading to the mature result that is still widely used today.
Unfortunately, since state functions are only fully known for the hydrogen atom, the quantum spin-orbit theory is only strictly appropriate for the description of D-line splitting in hydrogen or a one electron cation. However, this type of fine structure is known to appear in the spectra of many other elemental gases most notably the famous yellow doublet lines that appear in the spectra of Na. This doublet is depicted schematically in Figure 1.
Figure 1: Schematic energy level diagram showing the two transitions that yield the yellow sodium doublet.
In Figure 1 the spectroscopic notation $nl_{j}$, is used where $n$ is the principle quantum number $l$ the angular momentum quantum number where $l=S,P,D,\cdots$ and $j$ is the total angular momentum number and $j=l\pm 1/2$. In fact, all of the alkali metals display a very similar type of fine structure. We label the energy difference between the bands as $\Delta E_{so}$.
A well known method for estimating $\Delta E_{so}$ for hydrogen is through Rayleigh-Schrödinger perturbation theory where the spin-orbit effect is accounted for by using a perturbing spin-orbit Hamiltonian. The spin-orbit Hamiltonian can be derived from classical considerations and adjusted to include relativistic effects. Then, through the perturbation correction to the energy to first order, a closed form expression can be obtained for $\Delta E_{so}$. However, this result is only valid for the one electron atom and as such predicts the splitting energy reasonably well for hydrogen but fails for all of the heavier alkali metals.
When taking this approach the final expression for $\Delta E_{so}$ can be manipulated so that it is written in terms of the classic Bohr energy formula. We show here that by replacing this Bohr energy with the first ionization potential of the element involved that the expression then gives a greatly improved estimated for the D level splitting energies for all of the first column elements from Li to Fr. The best result was for Rb where the estimate for $\Delta E_{so}$ was within 3.2% of the experimentally reported value.
We then show that the original Bohr energy can be used to compute the correct D line splitting energy if it is adjusted with an appropriate screening constant. The well known Slater scheme for arriving at such a constant was considered but screening constants computed using this method did not yield accurate results. A method for generating the correct screening constants is discussed and utlilized here. Further, it is shown that these improved screening constants are a function of the atomic number with constant coefficients that can be estimated by the cation and atomic radii of the alkali metals.
The required ionization potentials were taken from experimental reports and computed using the General Atomic and Molecular Electronic Structure System (GAMESS) [2] at the density functional theory (DFT) level. All data and parameter values used to generate results mentioned in this report are listed in tables.
In the next section we briefly outline Rayleigh-Schrödinger first order perturbation theory and then, in greater detail, review the derivation of the spin-orbit Hamiltonian.
2 Theory
Rayleigh-Schrödinger perturbation theory can be used to obtain approximate solutions to a Schrödinger equation for cases where the differential equation cannot be resolved analytically by known techniques. To utilize this approach, it must be that the Hamiltonian can be divided into a two term Hamiltonian. One term, the unperturbed Hamiltonian, for which the analytic solution to the Schrödinger equation is known the other, the perturbing Hamiltonian, which then makes the system intractable. It is however a requirement that the effects of the perturbing Hamiltonian be small when compared to the non-perturbing part.
So, to deal with spin-orbit splitting in the hydrogen atom as a perturbation, the new Hamiltonian is then written in terms of the unperturbed, non-relativistic Hamiltonian for the hydrogen atom, $\hat{H}^{o}$, and a new perturbed Hamiltonian, $\hat{H}^{\prime}$, multiplied by a smallness factor $\lambda$. That is
$\hat{H}=\hat{H}^{o}+\lambda\hat{H}^{\prime}~{}.$ (1)
This method yields an expression for the new system energy eigenvalues, $E^{\prime}_{n}$ that are shifted somewhat due to the perturbation. The complete derivation for this result has been covered in detail previously and will not be fully reviewed here [3, 4, 5]. The final result to first order is
$E^{\prime}_{n}=E_{n}+\langle\Psi_{n}^{o}|\hat{H}_{so}|\Psi_{n}^{o}\rangle~{},$ (2)
where $E_{n}$ are the unperturbed Bohr energies, $\Psi_{n}^{o}$ the eigenfunctions of the unperturbed Hamiltonian and $\hat{H}_{so}$ the spin-orbit Hamiltonian where $\hat{H}_{so}=\lambda\hat{H}^{\prime}$.
We now require an expression for the spin-orbit Hamiltonian, $\hat{H}_{so}$. This begins by writing the potential energy for an electron orbiting a nucleus of $Z$ protons. Letting the rest frame be at the electron, the electron sees an orbiting nucleus of charge $Ze$ where $e$ is fundamental charge. Therefore, there is a magnetic field of magnitude $B$ at the location of the electron. From classical electrodynamics, $\hat{H}_{so}$, is then just the potential of the electron with magnetic moment $\vec{\mu}$ in the field of the nucleus given as
$\hat{H}_{so}=-\vec{\mu}\cdot\vec{B}~{}.$ (3)
Using the Biot-Savart law one can write $\vec{B}$ in terms of a current and then a magnetic dipole that can be given in terms of the angular momentum $\vec{L}$ for the orbit of the nucleus. The magnetic dipole moment of the electron is written in terms of its spin angular momentum $\vec{S}$. However, since the rest frame of the electron is not an inertial rest frame we must add a factor of $1/2$ to account for Thomas procession [6] so that
$H_{so}=\frac{Ze^{2}}{8\pi\epsilon_{0}}\frac{1}{m^{2}c^{2}r^{3}}\vec{S}\cdot\vec{L}~{}.$ (4)
This spin-orbit Hamiltonian involves the vectors $\vec{S}$ and $\vec{L}$ but we require an operator form. Therefore, we create a new vector, $\vec{J}$, called the total angular momentum where
$\vec{J}=\vec{S}+\vec{L}~{}.$ (5)
Using this relationship for the total angular momentum we can arrive at an operator form for $\vec{S}\cdot\vec{L}$. By evaluating the square of $\vec{S}+\vec{L}$, re-arranging and using Eq. (5), we arrive at
$\vec{S}\cdot\vec{L}=\frac{1}{2}[\hat{J}^{2}-\hat{L}^{2}-\hat{S}^{2}]~{}.$ (6)
Using Eq. (6) in Eq. (4) the spin-orbit Hamiltonian can be written as:
$H_{so}=\frac{Ze^{2}}{16\pi\epsilon_{0}}\frac{1}{m^{2}c^{2}r^{3}}[\hat{J}^{2}-\hat{L}^{2}-\hat{S}^{2}]~{}.$ (7)
Letting the first order correction energy be $E_{n}^{1}=\langle\Psi_{n}^{o}|\hat{H}_{so}|\Psi_{n}^{o}\rangle$ and then using $\hat{H}_{so}$ from Eq. (7) yields
$E_{n}^{1}=\frac{Ze^{2}}{16\pi\epsilon_{0}}\frac{1}{m^{2}c^{2}}\langle\psi_{n}|\frac{\hat{J}^{2}-\hat{L}^{2}-\hat{S}^{2}}{r^{3}}|\psi_{n}\rangle~{}.$ (8)
The required integration over $r^{-3}$ can be accomplished using a recursion formula and a generating function for the associated Laguerre polynomials [7]. Additionally, a method for computing this expectation value by making use of the Hellmann-Feynman theorem, is given by del Rio [8]. The eigenvalues are known for the operators $\hat{J}^{2}$, $\hat{L}^{2}$, and $\hat{S}^{2}$. Finally, after some rearranging, we can write the final result in terms of the Bohr energy and the fine structure constant $\alpha\approx 1/137$.
$E_{n}^{1}=\frac{Z^{2}|E_{n}|\alpha^{2}}{2n}\frac{[j(j+1)-l(l+1)-\frac{3}{4}]}{l(l+\frac{1}{2})(l+1)}~{}.$ (9)
Therefore, the energy of the D-line splitting, $\Delta E_{so}$, can be given by
$\Delta E_{so}=E_{n}^{1}\left(j=\frac{3}{2}\right)-E_{n}^{1}\left(j=\frac{1}{2}\right)~{}.$ (10)
Using Eq. (9) in this simply leads to
$\Delta E_{so}=\frac{Z^{2}(E_{n})\alpha^{2}}{2n}~{}~{}~{}\text{eV}~{}.$ (11)
3 Analysis
The result obtained in the previous section is used here to estimate $\Delta E_{so}$ for the alkali metals from Li to Fr. Obviously, we expect that when applying Eq. (9) which involves the one electron Bohr energy, the result will not be satisfactory. These calculations are carried out and compared with the experimental values for the D line splitting energies and listed in Table 1.
Table 1: D-line Energy Splitting in $10^{-3}$ eV
Atom    Experimental Values References    Eq. (10), Bohr $E_{n}$    $n$
Hydrogen    $4.39023\times 10^{-2}$  [9]    4.528 ×$\times 10^{-2}$    2
Lithium    $4.15641\times 10^{-2}$  [10]    3.668    2
Sodium    $2.133$  [11, 12]    196.0    3
Potassium    $7.159$  [13]    737.0    4
Rubidium    $29.486$  [14, 15, 16]    5429.0    5
Cesium    $68.756$  [17, 18, 19]    15340    6
Francium    —–    60481    7
It is clear from these data that when using the Bohr energies in Eq. (9), with $n$ as listed in Table 1 and $Z$ set to the atomic number of the element, the computed value for $\Delta E_{so}$ significantly overestimates the experimental value for all the alkali metals.
We propose here that an improved form for Eq. (9) can be obtained by replacing the Bohr energy, $E_{n}$, with the first ionization potential for the element involved, which we label as $\Delta E_{ion}$. Experimental values for first ionization potentials for the alkali metals have been tabulated [20]. To verify these values, first ionization potentials were also computed using the GAMESS software package [2] at the DFT/B3LYP level using the MIDI basis set for Li - Cs and the CRENBL effective core potential basis set for Fr [21]. These ionization potentials are listed in Table 2.
Table 2: First ionization potentials for hydrogen and the alkali metals in eV. From experiment and theory.
Atom    Exp. [20]    Computed [2]
Hydrogen    13.59844    13.52349
Lithium    5.39172    5.51992
Sodium    5.13908    5.21234
Potassium    4.34066    4.39092
Rubidium    4.17713    4.15699
Cesium    3.8939    3.76532
Francium    4.0727    1.14500
We now used the experimental ionization potentials in place of $E_{n}$ in Eq. (9) and re-computed values for $\Delta E_{so}$. These results are listed in Table 3. Obviously, there is a significant improvement over the calculations using the Bohr energy. On comparing these results to the experimental values in Table 1, we see that the best result was for Rb were there is an approximately 3.2% error between the theoretical result and the known experimental D-line splitting energy.
Table 3: D-line Energy Splitting values in $10^{-3}$ eV, from experiment and calculated using Eqs. (9) and (10) using experimental ionization potentials from Table 2 in place of Bohr energies. Far right column gives D-line splitting values using the Slater screening constants in Eq. (12).
Atom    Experimental Values References    Eq. (11)    Eq. (12) $S_{sl}$
Lithium    $4.15641\times 10^{-2}$  [10]    0.646    0.129
Sodium    $2.133$  [11, 12]    5.519    0.314
Potassium    $7.159$  [13]    10.43    0.166
Rubidium    $29.486$  [14, 15, 16]    30.452    1.061
Cesium    $68.756$  [17, 18, 19]    52.271    0.916
Francium    —–    117.315    0.840
4 Discussion
In the previous section we demonstrated how a known theory, used to estimated the D line splitting energy in the hydrogen atom, can be modified so as to yield estimates for this value for all of the alkali metals. The modification was the replacing of the Bohr energy in the final formula with the first ionization potential for the element involved. The energy values computed using this modified formula were in far greater agreement with the experimental values than those computed when using the Bohr energy. Clearly, the use of the ionization potential in the place of the Bohr energy helps correct the result by implicitly including screening effects in the multi-electron atoms.
Upon comparing our results for the D line splitting energy in Table 3 with the experimental values it can be seen that this scheme still overestimate the values for the elements lighter than Rb but underestimates the known value for Cs and likely also for Fr. That is, using an ionization potential to replace the Bohr energy in Eq. (9) underestimates screening for the lighter atoms but over estimates it for the heavier. We suspect that second order corrections to the energy will be too small to account for the remaining discrepancy and rather seek to find a corrected screening model for the Bohr energy that might be introduced in Eq. (9).
Many semi-empirical schemes have been developed for using the hydrogenic model to estimate total energies, polarizabilities and ionization potentials for multi-electron atoms by replacing $Z$ by an effective $Z$, $Z_{eff}$, where $Z_{eff}=Z-S$ [22, 23, 24, 25, 26]. Here $S$ is a screening constant. Also, the principle quantum number can be replaced with an effective principle quantum number $n^{*}$.
One of the most successful of these schemes is due to Slater [24]. Using Slater’s scheme, the screening constants for the outer $s$ electron in the group one atoms considered here were computed and labeled as $S_{sl}$. When directly using these values along with the corresponding values for $n^{*}$ [27] in Eq. (9), we find that the values for $\Delta E_{so}$ are off, often by an order of magnitude. These values are also listed in Table 3.
The Slater scheme can be used to generate accurate values for first ionization potentials, most notably, for elements in the first four rows of the periodic table. This is accomplished by finding the screening constant for the outer electron in the neutral atom and then again for the outer electron in the cation. Then, the total energy for each can be computed and the difference between these found. However, since the experimental values for the first ionization potentials fail to accurately describe the D level splitting energy for all of the alkali metals it must be that the Slater ionization potentials nor the Slater hydrogenic screening constants serve to yield accurate estimates for $\Delta E_{so}$ through Eq. (9). We therefore seek a screening constant that would be effective when using the Eq. (9) to describe the D line splitting in the alkali metals. It is a straight forward thing to use Eq. (11), and the known values for $\Delta E_{so}$, to compute an estimate for $S$ for each of the elements considered here. Eq. (11) adjusted to allowing for a screening constant is simply
$\Delta E_{so}=\frac{(Z-S)^{4}(13.6)\alpha^{2}}{2(n^{*})^{3}}~{}~{}~{}\text{eV}~{},$ (12)
where we let $n^{*}$ take on the Slater values [27]. Solving this for $S$ we get
$S=Z-\left[\frac{2(n^{*})^{3}\Delta E_{so}}{(13.6)\alpha^{2}}\right]^{1/4}~{},$ (13)
where $\Delta E_{so}$ is in eV. Using the experimental values for $\Delta E_{so}$, $S$ values for the alkali metals are computed using Eq. (13) and listed in Table 4 along with those generated using the Slater scheme.
Table 4: Screening constants, $S$, computed by using Eq. (13) and from using the Slater scheme for the outer $s$ electron, $S_{sl}$ [27]. ${}^{1}$stimated by the authors.
Atom    $S$ Eq. (14) $S_{sl}$ [27]    $Z$    $n^{*}$[27]
Lithium    2.02    1.7     3     2
Sodium    7.45    8.8     11     3
Potassium    13.37    16.8     19     3.7
Rubidium    28.50    33.3     37     4.0
Cesium    44.11    51.3     55     4.2
[…]
Obviously, the screening constants computed with Eq. (13) are a function of $Z$. A plot of these versus $Z$ is shown in Figure 2. We find that $S(Z)$ obeys a quadratic law given by
$S=a+bZ+cZ^{2}~{}.$ (14)
A curve of this type is fit to these data and shown in Figure 2. The fit was excellent having an R-squared value of 0.999.
Figure 2: Plot of $S$ values, computed using Eq. (13), vs. $Z$. The fit parameters from Eq. (14) were $a$ = -0.235754, $b$ = 0.686758 and $c$ = 0.00220759.
In spite of this excellent quadratic fit, the screening constants defined by Eq. (14) are very nearly linear with respect to $Z$.
We can find an estimate for the parameters $a$ and $b$ by considering the classical model for a positive point charge of magnitude $Ze$ in the center of a sphere of radius $R$ holding uniform negative charge density of total charge $-(Z-1)e$, where we neglect the charge of the outer $s$ electron. From Gauss’s law we get for the magnitude of the electric field, $\mathcal{E}$, at the position $r$ of the outer electron for $r<R$,
$\mathcal{E}=\frac{Ze-(Z-1)e\frac{r^{3}}{R^{3}}}{4\pi\epsilon_{o}r^{2}}~{}.$ (15)
With a little manipulation, one finds that the numerator of Eq. (15) can be written as
$e\left[Z-\left(Z\frac{r^{3}}{R^{3}}-\frac{r^{3}}{R^{3}}\right)\right]~{},$ (16)
so that we identify the screening constant as
$S=\left(\frac{r}{R}\right)^{3}(Z-1)~{}.$ (17)
To approximate the radii in Eq. (17) we let $r$ be the radius of the cation while $R$ is the atomic radius. This radii ratio varies from around 0.60 for Li to about 0.68 for Cs [28]. These values are consistent with the curve fit value for $b$. When this ratio is cubed, as in Eq. (17), the ratio ranges from 0.20 for Li to 0.32 for Cs. These numbers are more similar to the curve fit values for $a$. Therefore an approximate formula for the screening constant can be given by
$S=\left(\frac{r}{R}\right)Z-\left(\frac{r}{R}\right)^{3}~{},$ (18)
where the ratio $r/R$ could be set to the mean value for the group. We are lead to speculate that the mild quadratic behavior in the screening constant with increasing $Z$, is due to the onset of relativistic effects of the inner cores electrons on going down the group one column.
On examining Table 4, one finds that the Slater screening constants underestimate the screening for Li and overestimate screening for the other alkali metals. This was the same trend mentioned earlier when considering the effect of using ionization potentials to compute D line splitting energies which are listed in Table 3. This difference in the intensity of screening between the two models can be attributed to the electron-relaxation effect [25] whereby during ionization the atom contracts and thus slightly lowers the required energy for ionization from what one might expect when using a rigid, time-independent model of the atom. Therefore, the screening constants used to predict ionization energies implicitly correct for this effect by inflating the screening constant. However, electron relaxation is not an effect one must account for when modeling the spin-orbit phenomenon and therefore the screening constants in this case need not be overestimated.
5 Conclusion
In this report we demonstrated how a well known result for the D line splitting energy in hydrogen can be modified and then used to estimate this values for all of the alkali metals. The alteration is in replacing the Bohr energy with the first ionization potential for the element involved. This outcome then lead us to find values for screening constants that could be used to describe this effect whilst using the hydrogenic model. The screening constant is found to be a mildly quadratic function of the atomic number. Further, they produce a smaller screening effect than the screening constants generated to predict ionization potentials, a discrepancy we attribute to the electron relaxation effect.
By relating the problem to a classical situation from electrostatics, we are able to write an approximate formula for the screening constant in terms of the ratio of cation and atomic radii. It is hoped that this result might find use in predicting fine structure splitting energies for other atomic groups from the periodic table.
6 Acknowledgements
The authors wish to thank the College of Arts and Sciences at The University of the South for funding support.