### v2

# Estimating the D-line Energy Splitting of Alkali Metals Using a Relativistically Corrected Screening Constant

Abstract
We report here on how a known method from standard perturbation theory for estimating the energy of the D-line splitting in hydrogen can be modified to effectively approximate this quantity for all of the alkali metals. The approach utilizes a Rayleigh-Schrödinger perturbation theory first order correction to the energy. The perturbing Hamiltonian is the standard relativistically corrected spin-orbit Hamiltonian. From this, one derives an energy difference between the doublet lines that is theoretically appropriate for any one electron atom. This energy difference is written in terms of the Bohr energy. The results are good for hydrogen but, as expected, are significantly off from the experimental values for the multi-electron alkali metals. We show here that by replacing the Bohr energy with a first ionization potential, greatly improved values for the D-line splitting energy can be obtained from the model. However, this approach overestimates the splitting energy for the light alkali metals and underestimates it for the heavier ones. The best result was for Rb where the estimate only varied from the experimentally reported value by 3.2%. Accurate results are finally obtained from the model for all of the alkali metals when the original Bohr energy is adjusted with an appropriate screening constant. Screening constants generated using the Slater scheme however, which yield accurate estimates for ionization potentials, do not give the correct results for the D-line splitting energies. By fitting our results to reported energy splitting values we find that the correct screening constant is a function of the atomic number. We show that this result can be reproduced via a simple model with a relativistic correction for the contraction of the radius of the inner shell electrons.
Keywords: spin-orbit effect; sodium doublet; rubidium doublet; perturbation theory; screening constant
1 Introduction
The principle line spectra of an excited hydrogen gas sample were described empirically by J. J. Balmer in 1885. However, soon thereafter, an additional splitting in some of these principle lines was observed, an effect referred to as fine structure splitting. About a decade later, Michelson and Morley discovered fine structure splitting in the spectrum of sodium [1]. By ${}^{\text{th}}$ century, the Balmer line spectra of hydrogen had been described theoretically by the Bohr formula where lines were taken to be the result of energy transitions between quantum states with principle quantum number $n=2$ and $n=3,4,5,\cdots$.
After the appearance of Bohr’s formula, a quantitative explanation for fine structure in hydrogen spectra would soon be developed. Since that time, a variety of complex splittings have been observed in excited gas spectra but we will restrict our discussion here to a particular class of fine structure. This type of splitting occurs when one principle line is separated into two closely spaced bands, the so called doublets, or D-lines.
After the discovery of electron spin, a model for coupling between spin and orbital angular momentum was developed to explain the origin of the doublet splitting in hydrogen. This theory was further refined in 1927 by L. H. Thomas who included relativistic effects thus leading to the mature result that is still widely used today.
Unfortunately, since state functions are only fully known for the hydrogen atom, the quantum spin- orbit theory is only strictly appropriate for the description of D-line splitting in hydrogen or a one electron cation. However, this type of fine structure is known to appear in the spectra of many other elemental gases, most notably the famous yellow doublet lines that appear in the spectra of Na. This process is depicted schematically in Figure 1.
Figure 1: Schematic energy level diagram showing the two transitions that yield the yellow sodium doublet.
In Figure 1 the spectroscopic notation $nl_{j}$, is used where $n$ is the principle quantum number, $l$ the angular momentum quantum number where $l=S,P,D,\cdots$ and $j$ is the total angular momentum number and $j=l\pm 1/2$. In fact, all of the alkali metals display a very similar type of fine structure. We label the energy difference between the bands as $\Delta E_{so}$. A well-known method for estimating $\Delta E_{so}$ for hydrogen is through Rayleigh-Schrödinger perturbation theory where the spin- orbit effect is accounted for by using a perturbing spin-orbit Hamiltonian. The spin-orbit Hamiltonian can be derived from classical considerations and adjusted to include relativistic effects. Then, through the perturbation correction to the energy to first order, a closed form expression can be obtained for $\Delta E_{so}$. However, this result is only valid for the one electron atom and as such predicts the experimental splitting energy reasonably well for hydrogen but fails to accurately predict the same quantity for the heavier alkali metals.
When taking this approach, the final expression for $\Delta E_{so}$ can be manipulated so that it is written in terms of the classic Bohr energy $E_{n}$ , the fine structure constant $\alpha$, the atomic number $Z$ and the principle quantum number $n$. We show here that by replacing this Bohr energy with the first ionization potential of the element involved, this expression for $\Delta E_{so}$ yields values for the D-level splitting energies that are in better agreement with experimental results for all of the first column elements from Li to Fr. The best result was for Rb where the estimate for $\Delta E_{so}$ was within 3.2% of the experimentally reported value.
We then show that this model can be improved, and used to compute an accurate value for the D- line splitting energy, if the Bohr energy is adjusted with an appropriate screening constant. The well- known Slater scheme for arriving at such a constant was considered but screening constants computed using this method did not yield accurate results. From an empirical fit, it is shown that the correct screening constant is a function of the atomic number. We then demonstrate how a simple screening constant model, involving a relativistic correction for inner shell contraction, can be developed. The final result has the screening constant as a function of the atomic number with constant coefficients that can be related to the ratios of cationic to atomic radii.
The required ionization potentials were taken from experimental reports and computed using the General Atomic and Molecular Electronic Structure System (GAMESS) [2] at the density functional theory (DFT) level. All data and parameter values used to generate results mentioned in this report are listed in tables.
In the next section we briefly outline Rayleigh-Schrödinger first order perturbation theory and then, in greater detail, review the derivation of the spin-orbit Hamiltonian.
2 Theory
Rayleigh-Schrödinger perturbation theory can be used to obtain approximate solutions to a Schrödinger equation for cases where the differential equation cannot be resolved analytically by known techniques. To utilize this approach, one typically divides the Hamiltonian into two terms. One term, the unperturbed Hamiltonian, for which the analytic solution to the Schrödinger equation is known and, the other, the perturbing Hamiltonian, which then makes the system intractable. It is however a requirement that the effects of the perturbing Hamiltonian be small when compared to the non- perturbing part.
So, to deal with spin-orbit splitting in the hydrogen atom as a perturbation, the new Hamiltonian is then written in terms of the unperturbed, non- relativistic Hamiltonian for the hydrogen atom, $\hat{H}^{o}$, , and a new perturbed Hamiltonian, $\hat{H}^{\prime}$, , multiplied by a smallness factor $\lambda$. That is
$\hat{H}=\hat{H}^{o}+\lambda\hat{H}^{\prime}~{}.$ (1)
This method yields an expression for the new system energy eigenvalues, $E^{\prime}_{n}$, that are shifted somewhat due to the perturbation. The complete derivation for this result has been covered in detail previously and will not be fully reviewed here [3, 4]. The final result to first order is
$E^{\prime}_{n}=E_{n}+\langle\Psi_{n}^{o}|\hat{H}_{so}|\Psi_{n}^{o}\rangle~{},$ (2)
where $E_{n}$ are the unperturbed Bohr energies, $\Psi_{n}^{o}$ the eigenfunctions of the unperturbed Hamiltonian and $\hat{H}_{so}$ the spin-orbit Hamiltonian where $\hat{H}_{so}=\lambda\hat{H}^{\prime}$.
We now require an expression for the spin-orbit Hamiltonian, $\hat{H}_{so}$. This begins by writing the potential energy for an electron orbiting a nucleus of $Z$ protons. Letting the rest frame be at the electron, the electron sees an orbiting nucleus of charge $Ze$ where $e$ is fundamental charge. Therefore, there is a magnetic field of magnitude $B$ at the location of the electron. From classical electrodynamics, $\hat{H}_{so}$, is then just the potential of the electron with magnetic moment $\vec{\mu}$ in the field of the nucleus given as
$\hat{H}_{so}=-\vec{\mu}\cdot\vec{B}~{}.$ (3)
Using the Biot-Savart law one can write $\vec{B}$ terms of the current, as seen by the electron, and then this can be further written in terms of the orbital angular momentum, $\vec{L}$. The magnetic dipole moment of the electron is then given in terms of its spin angular momentum $\vec{S}$. However, since the rest frame of the electron is not an inertial rest frame, we must add a factor of $1/2$ to account for Thomas procession so that
$H_{so}=\frac{Ze^{2}}{8\pi\epsilon_{0}}\left(\frac{1}{m^{2}c^{2}r^{3}}\right)\vec{S}\cdot\vec{L}~{}.$ (4)
This spin-orbit Hamiltonian involves the vectors $\vec{S}$ and $\vec{L}$ but we require an operator form. Therefore, we create a new vector, $\vec{J}$, called the total angular momentum where
$\vec{J}=\vec{S}+\vec{L}~{}.$ (5)
Using this relationship for the total angular momentum we can arrive at an operator form for $\vec{S}\cdot\vec{L}$. By evaluating the square of $\vec{S}+\vec{L}$, re-arranging and using Eq. (5), we arrive at
$\vec{S}\cdot\vec{L}=\frac{1}{2}[\hat{J}^{2}-\hat{L}^{2}-\hat{S}^{2}]~{}.$ (6)
Using Eq. (6) in Eq. (4) the spin-orbit Hamiltonian can be written as:
$H_{so}=\frac{Ze^{2}}{16\pi\epsilon_{0}}\left(\frac{1}{m^{2}c^{2}r^{3}}\right)[\hat{J}^{2}-\hat{L}^{2}-\hat{S}^{2}]~{}.$ (7)
Letting the first order correction energy be $E_{n}^{1}=\langle\Psi_{n}^{o}|\hat{H}_{so}|\Psi_{n}^{o}\rangle$ and then using $\hat{H}_{so}$ from Eq. (7) yields
$E_{n}^{1}=\frac{Ze^{2}}{16\pi\epsilon_{0}}\left(\frac{1}{m^{2}c^{2}}\right)\langle\psi_{n}|\frac{\hat{J}^{2}-\hat{L}^{2}-\hat{S}^{2}}{r^{3}}|\psi_{n}\rangle~{}.$ (8)
The required integration over $r^{-3}$ can be accomplished using a recursion formula and a generating function for the associated Laguerre polynomials [5]. The eigenvalues are known for the operators $\hat{J}^{2}$, $\hat{L}^{2}$, and $\hat{S}^{2}$. Finally, after some rearranging, we can write the final result in terms of the Bohr energy and the fine structure constant $\alpha\approx 1/137$.
$E_{n}^{1}=\frac{Z^{2}|E_{n}|\alpha^{2}}{2n}\frac{[j(j+1)-l(l+1)-\frac{3}{4}]}{l(l+\frac{1}{2})(l+1)}~{}.$ (9)
Therefore, the energy of the D-line splitting, $\Delta E_{so}$, can be given by
$\Delta E_{so}=E_{n}^{1}\left(j=\frac{3}{2}\right)-E_{n}^{1}\left(j=\frac{1}{2}\right)~{}.$ (10)
Using Eq. (9) in this simply leads to
$\Delta E_{so}=\frac{Z^{2}|E_{n}|\alpha^{2}}{2n}~{}.$ (11)
3 Analysis
The result obtained in the previous section is used here to estimate $\Delta E_{so}$ for the alkali metals from Li to Fr. Obviously, we expect that when applying Eq. (11) which involves the one electron Bohr energy, the result will not be satisfactory. These calculations were carried out and compared with the experimental values for the D-line splitting energies and listed in Table 1.
Table 1: D-line Energy Splitting in $10^{-3}$ eV
Atom      Experimental Values References      Eq. (10), Bohr $E_{n}$      $n$
Hydrogen      $4.39023\times 10^{-2}$ [6]      4.528 ×$\times 10^{-2}$      2
Lithium      $4.15641\times 10^{-2}$ [7]      3.668      2
Sodium      $2.133$ [8, 9]      196.0      3
Potassium      $7.159$ [10]      737.0      4
Rubidium      $29.486$ [11, 12, 13]      5429.0      5
Cesium      $68.756$ [14, 15, 16]      15340      6
Francium      —–      60481      7
It is clear from these data that when using the Bohr energies in Eq. (9), with $n$ as listed in Table 1 and $Z$ set to the atomic number of the element, the computed value for $\Delta E_{so}$ significantly overestimates the experimental value for all the alkali metals.
We propose here that an improved form for Eq. (9) can be obtained by replacing the Bohr energy, $E_{n}$, with the first ionization potential for the element involved, which we label as $\Delta E_{ion}$. Experimental values for first ionization potentials for the alkali metals have been tabulated [17]. To verify these values, first ionization potentials were also computed using the GAMESS software package [2] at the DFT/B3LYP level using the MIDI basis set for Li - Cs and the CRENBL effective core potential basis set for Fr [18]. These ionization potentials are listed in Table 2.
Table 2: First ionization potentials for hydrogen and the alkali metals in eV. From experiment and theory.
Atom      Exp. [17]      Computed [2]
Hydrogen      13.59844      13.52349
Lithium      5.39172      5.51992
Sodium      5.13908      5.21234
Potassium      4.34066      4.39092
Rubidium      4.17713      4.15699
Cesium      3.8939      3.76532
Francium      4.0727      1.14500
We now used the experimental ionization potentials in place of $E_{n}$ in Eq. (9) and re-computed values for $\Delta E_{so}$. These results are listed in Table 3. Obviously, there is a significant improvement over the calculations using the Bohr energy. On comparing these results to the experimental values in Table 1, we see that the best result was for Rb were there is an approximately 3.2% error between the theoretical result and the known experimental D-line splitting energy.
Table 3: D-line Energy Splitting values in $10^{-3}$ eV, from experiment and calculated using Eqs. (9) and (10) using experimental ionization potentials from Table 2 in place of Bohr energies. Far right column gives D-line splitting values using the Slater screening constants in Eq. (12).
Atom      Experimental Values References      Eq. (11)      Eq. (12) $S_{sl}$
Lithium      $4.15641\times 10^{-2}$ [7]      0.646      0.129
Sodium      $2.133$ [8, 9]      5.519      0.314
Potassium      $7.159$ [10]      10.43      0.166
Rubidium      $29.486$ [11, 12, 13]      30.452      1.061
Cesium      $68.756$ [14, 15, 16]      52.271      0.916
Francium      —–      117.315      0.840
4 Discussion
In the previous section we demonstrated how a known theory, used to estimate the D-line splitting energy in the hydrogen atom, can be modified so as to yield estimates for these values for all of the alkali metals. The modification was the replacing of the Bohr energy in the final formula with the first ionization potential for the element involved. The energy values computed using this modified formula were in far greater agreement with the experimental values than those computed when using the Bohr energy. Clearly, the use of the ionization potential in the place of the Bohr energy helps correct the result by implicitly including screening effects in the multi-electron atoms.
Upon comparing our results for the D line splitting energy in Table 3 with the experimen- tal values it can be seen that this scheme still overestimate the values for the elements lighter than Rb but underestimates the known value for Cs and likely also for Fr. That is, using an ionization potential to replace the Bohr energy in Eq. (9) underestimates screening for the lighter atoms but over estimates it for the heavier. We suspect that second order corrections to the energy will be too small to account for the remaining discrepancy and rather seek to find a corrected screening model for the Bohr energy that might be introduced in Eq. (9).
Many semi-empirical schemes have been developed for using the hydrogenic model to estimate total energies, polarizabilities and ionization potentials for multi-electron atoms by replacing $Z$ by an effective $Z$, $Z_{eff}$, where $Z_{eff}=Z-S$ [19, 20, 21, 22]. Here $S$ is a screening constant. Also, the principle quantum number can be replaced with an effective principle quantum number $n^{*}$.
A well-known method for computing the hydrogenic screening constant was developed, in the early twentieth century, by J. C. Slater [19]. Using this prescription, we computed the screening constant for the outer $s$ electron in the group one atoms and labeled these as $S_{sl}$ These screening constant values are listed in Table 3. When directly using these values in $Z_{eff}$ , along with the corresponding values for $n^{*}$ listed in Reference [22], in Eq. (9), we find that the values for $\Delta E_{so}$ are off, often by an order of magnitude. These values are also listed in Table 3.
The Slater scheme can be used to generate accurate values for first ionization potentials, most notably, for elements in the first four rows of the periodic table. This is accomplished by finding the screening constant for the outer electron in the neutral atom and then again for the outer electron in the cation. Then, the total energy for each can be computed and the difference between these found. However, since the experimental values for the first ionization potentials, when used in Eq. (11) in place of the Bohr energy, fail to yield accurate values for the D-line splitting energy for all of the alkali metals it is no surprise that Slater ionization potentials, nor the Slater hydrogenic screening constants, would serve to yield accurate estimates for $\Delta E_{so}$ either. We therefore seek a screening constant that would be effective when using Eq. (11) to compute the D-line splitting energies in the alkali metals.
First we use the experimental data for $\Delta E_{so}$ to find an empirical relation for S. It is a straight forward thing to use Eq. (11), and the known values for $\Delta E_{so}$ , to compute an estimate for S for each of the elements considered here. Eq. (11) adjusted to allowing for a screening constant is simply
$\Delta E_{so}=\frac{(Z-S)^{4}(13.6)\alpha^{2}}{2(n^{*})^{3}}~{}~{}~{}\text{eV}~{},$ (12)
where we let $n^{*}$ take on the Slater values [22] and $\Delta E_{so}$ has units of electron-volt (eV). Solving this for $S$ we get
$S=Z-\left[\frac{2(n^{*})^{3}\Delta E_{so}}{(13.6)\alpha^{2}}\right]^{1/4}~{},$ (13)
Using the experimental values for $\Delta E_{so}$ , screening constants for the alkali metals are computed using Eq. (13) and listed in Table 4 along with those generated using the Slater scheme.
Table 4: Screening constants, $S$, computed by using Eq. (13) and $S_{sl}$ computed using the Slater scheme for the outer $s$ electron. Effective principle quantum numbers were taken from [22] except in the case of Francium.
Atom      $S$ Eq. (14) $S_{sl}$ [22]      $Z$      $n^{*}$[22]
Lithium      2.02      1.7       3       2
Sodium      7.45      8.8       11       3
Potassium      13.37      16.8       19       3.7
Rubidium      28.50      33.3       37       4.0
Cesium      44.11      51.3       55       4.2
Francium      76.16      83.3       87       4.31
Obviously, the screening constants computed with Eq. (13) are a function of $Z$. A plot of these versus $Z$ is shown in Figure 2. We find that $S(Z)$ obeys a quadratic law given by
$S=a+bZ+cZ^{2}~{}.$ (14)
A curve of this type is fit to these data and shown in Figure 2. The fit was excellent having an R-squared value of 0.999.
Figure 2: Plot of $S$ values, computed using Eq. (13), vs. $Z$. The fit parameters from Eq. (14) were $a$ = -0.235754, $b$ = 0.686758 and $c$ = 0.00220759.
We develop a simple model for $S$ by considering the classical model for the electric field magnitude, $\mathcal{E}$ at the outer $s$ electron. Consider a Gaussian spherical surface of radius $r$ which encloses the charge of the nucleus, $Ze$ and some fraction of the charge, $-(Z-1)e$. due to the electron cloud excluding the outer electron. Quantum mechanics predicts that the inner electrons may spend part of the time at greater distance from the nucleus that the outer $s$ electron so that, when averaged over time, some fraction of the electron cloud is enclosed within the Gaussian surface at the location $r$, of the outer $s$ electron. In Gauss’s law this gives
$\mathcal{E}=\frac{Ze-(Z-1)e\frac{r^{3}}{R^{3}}}{4\pi\epsilon_{o}r^{2}}~{},$ (15)
where $r<R$. With a little manipulation, one finds that the numerator of the right side of Eq. (15) can be written as
$e\left[Z-\left(Z\frac{r^{3}}{R^{3}}-\frac{r^{3}}{R^{3}}\right)\right]~{},$ (16)
so that we identify the screening constant as
$S=\left(\frac{r}{R}\right)^{3}(Z-1)~{}.$ (17)
The exact value to use for the radius $r$ is unknown but, we expect that it will vary with $Z$ due to relativistic effects. It is known that the velocity of inner shell electrons, being directly proportional to $Z$, can approach a sizable fraction of the speed of light which in turn leads to mass enhancement which contracts the orbital and causes a higher screening of the nuclear charge as $Z$ increases [23]. Therefore we propose that, due to such effects effects, the value for $r$ contracts on going down the group one column. This can be modeled by an expression of the form [24, 25].
$r=r_{o}\sqrt{1-\left(\frac{Z}{nc}\right)^{2}}~{}.$ (18)
Here $n$ is a quantum number and $c$ the speed of light. Clearly $Z/(nc)<<1$ so that on cubing both sides of Eq. (18) it can be expanded to arbitrary order in a Maclaurin series as
$r^{3}={r_{o}}^{3}\left(1-\frac{3}{2}\left(\frac{Z}{nc}\right)^{2}+\frac{3}{8}\left(\frac{Z}{nc}\right)^{4}+\frac{1}{16}\left(\frac{Z}{nc}\right)^{6}+\frac{3}{128}\left(\frac{Z}{nc}\right)^{8}+\mathcal{O}^{8}\right)~{}.$ (19)
Using Eq. (19) for $r$ in Eq. (17), then rearranging and factoring yields
$S=\left(\frac{r_{o}}{R}\right)^{3}\left[Z-1+Z^{2}\left(\frac{3}{2(nc)^{2}}-\frac{3Z}{2(nc)^{2}}-\frac{3Z^{2}}{8(nc)^{4}}+\frac{3Z^{3}}{8(nc)^{4}}-\frac{Z^{4}}{16(nc)^{6}}+\mathcal{O}^{7}\right)\right]~{}.$ (20)
Now, in this case $3\leq Z\leq 87$ and as $nc>>Z$ the terms in parenthesis of Eq. (20) are certainly small and getting progressively smaller so that we set the sequence in parenthesis to be the constant $\gamma$. Therefore, the model gives as a final expression for the screening constant
$S=-\left(\frac{r_{o}}{R}\right)^{3}+\left(\frac{r_{o}}{R}\right)^{3}Z+\left(\frac{r_{o}}{R}\right)^{3}\gamma Z^{2}~{},$ (21)
which is in agreement with the empirical result given in Eq. (14).
The precise value to use for the constant $r_{o}$ is not known to the authors. However, as a first estimate for the ratio $r_{o}/R$ we let $r_{o}$ be the cation radius and, as mentioned previously, $R$ is the covalent radius. From tabulated data [26], one finds that $(r_{o}/R)^{3}$ ranges from 0.208 for Li to 0.319 for Cs. On comparing Eqs. (14) and (21) we have $a=b=(ro/R)^{3}$. The values for the constants $a$ and $b$ from the curve fit of Figure 2 are $a=0.236$ and $b=0.687$. As the agreement between $(r_{o}/R)^{3}$ and $a$ is excellent for Li and, within an order of magnitude for $b$, one must assume we are on the right track and we therefore suggest that the cube of the cation to covalent radius would be a reasonable choice for $a$ and $b$ when using Eq. (21) to estimate a screening constant. Further research is necessary to clarify the exact nature of $r_{o}$.
On examining Table 4, one finds that the Slater screening constants underestimate the screening for Li and overestimate screening for the other alkali metals. This was the same trend mentioned earlier when considering the effect of using ionization potentials to compute D-line splitting energies which are listed in Table 3. This difference in the intensity of screening between the two models can be attributed to the electron-relaxation effect [20] whereby during ionization the atom contracts and thus slightly lowers the required energy for ionization from what one might expect when using a rigid, time-independent model of the atom. Therefore, the screening constants used to predict ionization energies implicitly correct for this effect by inflating the screening constant. However, electron relaxation is not an effect one must account for when modeling the spin-orbit phenomenon and therefore the screening constants in this case need not be overestimated.
5 Conclusion
In this report we demonstrated how a well known result for the D-line splitting energy in hydrogen can be modified and then used to estimate this value for all of the alkali metals. The alteration is in adjusting the Bohr energy with an appropriate screening constant which is a function of the atomic number. Through an empirical fit the screening constant is shown to be a mildly quadratic function of the atomic number. Using a simple model, we show how this form of the screening constant can be derived from first principles with a relativistic correction. It is hoped that this result might find use in predicting fine structure splitting energies for other atomic groups from the periodic table.
6 Acknowledgements
The authors wish to thank the College of Arts and Sciences at The University of the South for funding support.